OpenStudy (raffle_snaffle):

In a class of 100 students, 40 are computer science majors, 36 are mechanical engineering majors, 18 are civil engineers and the rest are general engineering majors. Assume students only have one major. If a student is chosen at random what is the probability they are: (Round all answers to two decimal places) a) a civil engineering major or mechanical engineering major? b) Suppose five students from the class are chosen at random. What is the probability none are mechanical engineering majors?

OpenStudy (raffle_snaffle):

@tkhunny

OpenStudy (raffle_snaffle):

For a is it P = (36+18)/100 = 0.54

OpenStudy (tkhunny):

We're just counting. (18+36)/100 -- Why is it harder than that? Second is a little trickier. Shall we approximate or produce the exact value?

OpenStudy (bahrom7893):

Probability that 5 chosen at random and non are mech eng: 1/(100-36) * 1/(99-36)... Am I right?

OpenStudy (bahrom7893):

Sorry it's been a while, wanted to post before seeing tk's solution

OpenStudy (raffle_snaffle):

LOL because I did this problem and I got it wrong... I added 6 the first time byt now I see why my other answer makes more sense.

OpenStudy (raffle_snaffle):

I don't know if you are right.

OpenStudy (raffle_snaffle):

I don't have the answers.

jimthompson5910 (jim_thompson5910):

`36 are mechanical engineering majors` so 100-36 = 63 are majored in something else x = 63 C 5 = number of ways to pick 5 non-mechanical engineering students y = 100 C 5 = number of ways to pick any 5 students the answer will be in the form x/y I'm using this combination formula (n C r) \[\Large _n C_r = \frac{n!}{r!*(n-r)!}\]

jimthompson5910 (jim_thompson5910):

sorry not 63, made a typo should be 64

OpenStudy (raffle_snaffle):

Where did you get that equation? *looking in text book*

OpenStudy (bahrom7893):

@jim_thompson5910 and sorry for turning your question into a discussion @raffle_snaffle , but how would I approach this if I didn't remember the formulas?

jimthompson5910 (jim_thompson5910):

look for the combinations or permutations section

OpenStudy (raffle_snaffle):

I don't know @bahrom7893

OpenStudy (bahrom7893):

(taking probability seminar in a week, need to review all these things myself and I have a terrible memory for formulas)

OpenStudy (raffle_snaffle):

I don't think we have covered that formula yet... Can I use Binomial distribution or Possion DIstrbution?

jimthompson5910 (jim_thompson5910):

think of 5 slots (A,B,C,D,E) if you have 100 people to choose from 100 choices for slot A 99 for slot B 98 for slot C 97 for slot D 96 for slot E multiply those choices out 100*99*98*97*96 = 9,034,502,400 9,034,502,400 is the number of ways to pick 5 out of a pool of 100 where order matters. This is equal to 100 P 5 since order doesn't matter, you divide by 5! = 120 9,034,502,400/120 = 75,287,520 this is exactly equal to 100 C 5

jimthompson5910 (jim_thompson5910):

I divided by 5! = 120 because there are 120 ways to arrange 5 items eg: ABCDE is the same as ABCED

OpenStudy (raffle_snaffle):

Okay @jim_thompson5910

OpenStudy (triciaal):

why wouldn't you use 64?

OpenStudy (bahrom7893):

Nvm I just saw your example

OpenStudy (raffle_snaffle):

I am going to ask my instructor next week since I have one more attempt for this problem. Thanks for the help everyone.